# Privacy-Aware Detection of Fake Identity Documents: Methodology, Benchmark, and Improved Detection Methods (FakeIDet2)

### v1

@@ 6.6 Cross-Database Scenario
This section evaluates the performance of FakeIDet2 trained through the selected configuration (i.e., Pseudo-Anonymized $\times64$ patch size) on different public databases in the literature. Concretely, we consider DLC-2021 [30], KID34K [27] and Benalcazar et al. synthetic database from [5]. For this evaluation, we sample around 10,000 images from KID34K, balancing across PAIs and ID owners, and roughly 30,000 images from several videos of DLC-2021, sampling in a similar way as in KID34K. DLC-2021 contains Spanish IDs, but they are from the first version, which are not seen during training FakeIDet2 as described in Sec. 5.1. Regarding the real IDs, as in those databases the “real” IDs were created by the researchers in laboratory conditions, i.e., they are not official IDs, we consider the aforementioned databases as attacks. Real IDs are extracted from the evaluation dataset of our FakeIDet2-db, similar to previous experiments, as they are official IDs. Our aim with this privacy-aware scenario is to assess the generalization ability of our proposed FakeIDet2 to unseen conditions, e.g., different PAIs, templates, devices, etc.
@@
Table 9 shows the performance of FakeIDet2 in terms of EER (%) per type of attack and database. In order to facilitate the collaboration between ID Holders and AI Researchers, we consider the privacy-aware scenario where FakeIDet2 is trained using the Pseudo-Anonymized $\times64$ patch size configuration. For DLC-2021 and KID34K we denote as “HQ-print” the types of print attacks which are proposed as “real” in each of the databases, and for Benalcazar et al. we denote “Synthetic” for the samples that have been created via generative methods trained from real samples. Additionally, in case a database does not contain a specific attack, it is denoted as “N/A”.
As can be seen in Table 9, the performance per PAI varies among databases. For instance, FakeIDet2 seems to perform worst in screen attacks from the KID34K database compared to the DLC-2021, with 13.41% vs. 5.02%, respectively. This could be motivated by the amount of devices which were used for screen attack creation in KID34K, which varied from smartphones, tablets, laptops, and LCD monitors. A similar trend can be observed for the HQ-print attacks, where KID34K samples seem to confuse FakeIDet2 more (18.26% vs. 10.28%). This may be induced due to the variability in terms of acquisition devices (12 in KID34K vs. 2 in DLC-2021), with different specifications that create different traces or fingerprints in the images taken. Regarding print attacks, we see that our model performs better on the KID34K dataset than in DLC-2021 (4.99% vs. 12.45%). This might be produced as KID34K follows a similar approach to ours, i.e., print attacks are created after printed and then laminated. Finally, for the synthetic attacks included in Benalcazar et al. database, FakeIDet2 is not able to generalize well to this unseen attack (39.41% EER). This result is expected due to the following reasons: i) the data (both real and fake ID samples) used to train FakeIDet2 is only based on physical acquisitions using smartphone cameras, not synthetic data, and ii) according to [5], those synthetic samples were generated using GANs trained with official Chilean ID samples, which means they share similar characteristics with real IDs that may be confusing FakeIDet2.
In conclusion, our proposed privacy-aware scenario in order to promote the collaboration between ID Holders and AI Researchers has provided very interesting results, despite of the fact that our proposed FakeIDet2 is only trained through the Pseudo-Anonymized $\times64$ patch size configuration, using the the development set of FakeIDet2-db. Under unseen physical attacks, FakeIDet2 is able to achieve 8.90% EER on DLC-2021 and 13.84% on KID34K, as can be seen in Table 9. We expect these results will further improve in real-world settings, as the ID Holders own large-scale ID databases.
@@ 7 Conclusion
In this article we have explored a novel framework for privacy-aware fake ID detection, which bridges the existing gap between IDs Holders (e.g., governments, police, banks, etc.), which provide digital services to citizens and own large-scale datasets of real IDs, and the AI Researchers, which have the experience and technology to develop fake ID detectors, but not the data.
In order to advance in the field, we have first introduced FakeIDet2-db, a database that complies with the proposed framework, which contains over 900K patches in different sizes ($\times128$ and $\times64$), extracted from a total of 2,000 images of pseudo- and fully-anonymized IDs. Regarding real and fake IDs, to the best of our knowledge this is the first database that contains: i) official real IDs captured under several conditions regarding camera sensor, illumination conditions and distance, and ii) physical composite attacks. In addition, our database considers a large variability during the acquisition, e.g., using three different smartphones, and changes in illumination and height conditions.
Along with FakeIDet2-db, we have presented FakeIDet2, a new method for privacy-aware fake ID detection that leverages patches from IDs and varying levels of anonymization to safeguard the sensitive information. We demonstrate that learning embeddings from patches extracted from real and fake IDs is a crucial step for effective fake ID detection. Considering the nature of the problem, especially for physical composite attacks, we have proposed a method that explicitly accounts for class imbalance, achieving superior performance compared to traditional approaches. Using patch-level embeddings, we have incorporated a learnable fusion module based on MHSA to model the relationships between patches within a single ID, producing a ID-level prediction indicating whether the submitted ID is real or fake. Additionally, we have evaluated the impact of different anonymization levels and patch sizes, showing that our FakeIDet2 remains competitive in privacy-aware scenarios, as compared to fully visible data settings (3.99% in non-anonymized vs. 8.64% in pseudo-anonymized).
Finally, we have also introduced an extensive and reproducible public benchmark, which considers in addition to our FakeIDet2-db, other public databases in the literature, containing unseen physical (KID34K [27], DLC-2021 [30]) and synthetic attacks (Benalcazar et al. [5]) to assess the performance on different data distributions that contain IDs from several demographic regions (i.e., Europe, Asia and South America).

### v2

@@ 6.6 Cross-Database Scenario
This section evaluates the performance of FakeIDet2 trained through the selected configuration (i.e., Pseudo-Anonymized $\times64$ patch size) on different public databases in the literature. Concretely, we consider DLC-2021 [30], KID34K [28] and Benalcazar et al. synthetic database from [5]. For this evaluation, we sample around 10,000 images from KID34K, balancing between PAIs and ID owners, and approximately 30,000 images from several DLC-2021 videos, sampling in a similar way as in KID34K. DLC-2021 contains Spanish IDs, but they are from the first version, which are not seen during training FakeIDet2 as described in Sec. 5.1. Regarding the real IDs, as in those databases the “real” IDs were created by the researchers under laboratory conditions, that is, they are not official IDs, we consider the aforementioned databases as attacks. Real IDs are extracted from the evaluation dataset of our FakeIDet2-db, similar to previous experiments, since they are official IDs. Our aim with this privacy-aware scenario is to assess the generalization ability of our proposed FakeIDet2 to unseen conditions, e.g., different PAIs, templates, devices, etc.
@@
Table 9 shows the performance of FakeIDet2 in terms of EER (%) per type of attack and database. In order to facilitate the collaboration between ID Holders and AI Researchers, we consider the privacy-aware scenario where FakeIDet2 is trained using the Pseudo-Anonymized $\times64$ patch size configuration. For DLC-2021 and KID34K we denote by “HQ-Print” the types of print attacks which are proposed as “real” in each of the databases, and for Benalcazar et al. we denote “Synthetic” the samples that have been created via generative methods trained from real samples. Furthermore, in case a database does not contain a specific attack, it is denoted as “N/A”.
As can be seen in Table 9, the performance per PAI varies between databases. For example, FakeIDet2 is worse in screen attacks from the KID34K database compared to the DLC-2021, with 13.41% vs. 5.02% EER, respectively. This could be motivated by the amount of devices that were used to create screen attacks in KID34K, ranging from smartphones, tablets, laptops, and LCD monitors. A similar trend can be observed for the HQ-Print attacks, where the KID34K samples seem to confuse FakeIDet2 more (18.26% vs 10.28% EER). This may be induced by the variability in terms of acquisition devices (12 in KID34K vs. 2 in DLC-2021), with different specifications that create different traces or fingerprints in the images taken. Regarding print attacks, we see that our model performs better on the KID34K dataset than in DLC-2021 (4.99% vs. 12.45%). This might be produced because KID34K follows a similar approach to ours, i.e., print attacks are created after printed and then laminated. Finally, for the synthetic attacks included in Benalcazar et al. database, FakeIDet2 is not able to generalize well to this unseen attack (39.41% EER). This result is expected due to the following reasons: i) the data (both real and fake ID samples) used to train FakeIDet2 are only based on physical acquisitions using smartphone cameras, not synthetic data, and ii) according to [5], those synthetic samples were generated using GANs trained with official Chilean ID samples, which means that they share similar characteristics with real IDs that may confuse FakeIDet2.
In conclusion, our proposed privacy-aware scenario in order to promote the collaboration between ID Holders and AI Researchers has provided very interesting results, despite the fact that our proposed FakeIDet2 is only trained through the Pseudo-Anonymized $\times64$ patch size configuration, using the development set of FakeIDet2-db. Under unseen physical attacks, FakeIDet2 is able to achieve 8.90% EER on DLC-2021 and 13.84% on KID34K, as can be seen in Table 9. We expect that these results can be enhanced in real-world settings, as the ID Holders own large-scale ID databases that can be exploited to improve the machine learning processes described here.
@@ 7 Conclusion
In this article, we have explored a novel framework for privacy-aware fake ID detection, which bridges the existing gap between IDs Holders (e.g., governments, police, banks, etc.), which provide digital services to citizens and own large-scale datasets of real IDs, and AI Researchers, which have the experience and technology to develop fake ID detectors, but not the data.
In order to advance in the field, we have first introduced FakeIDet2-db, a database that complies with the proposed framework, which contains more than 900K patches in different sizes ($\times128$ and $\times64$), extracted from a total of 2,000 images of pseudo-anonymized and fully-anonymized IDs. Regarding real and fake IDs, to our knowledge, this is the first database that contains: i) official real IDs captured under several conditions regarding camera sensor, illumination conditions and distance, and ii) physical composite attacks. In addition, our database considers a large variability during the acquisition, e.g., using three different smartphones, and changes in illumination and height conditions.
Together with FakeIDet2-db, we have presented FakeIDet2, a new method for the detection of fake ID that is sensitive to privacy and utilizes patches from IDs and varying levels of anonymization to protect sensitive information. We show that learning embeddings from patches extracted from real and fake IDs is a crucial step for effective fake ID detection. Considering the nature of the problem, especially for physical composite attacks, we have proposed a method that explicitly accounts for class imbalance using time-adaptive class weights [14] during learning and have compared it to other popular loss functions [23], achieving superior performance compared to traditional approaches (CosFace, ArcFace, and AdaFace). Using patch-level embeddings, we have incorporated a learnable fusion module based on multi-head self-attention [4] to assign different weights to different patches within a single ID, producing an ID-level prediction that indicates whether the submitted ID is real or fake. Furthermore, we have evaluated the impact of different anonymization levels and patch sizes, showing that our FakeIDet2 remains competitive in privacy-aware scenarios, as compared to fully visible data settings (3.99% in non-anonymized vs. 8.64% in pseudo-anonymized).
Finally, we have also introduced an extensive and reproducible public benchmark, which considers, in addition to our FakeIDet2-db, other public databases in the literature, containing unseen physical (KID34K [28], DLC-2021 [30]) and synthetic attacks (Benalcazar et al. [5]) to assess performance in different data distributions that contain IDs from several demographic regions (i.e., Europe, Asia, and South America).
@@
In future work, we will explore other foundation models [27] and other learning architectures based on the patch-based principles shown in Fig. 6 and other non-patch and inpainting detection methods [8] following ongoing benchmarks and competitions such as DeepID [^11] https://deepid-iccv.github.io/. Improving the accuracy of FakeIDet2 by exploiting the spatial context of patches (e.g., using spatial attention [10] and local quality measures [31]) and a full layout analysis [19] is also in our plans.